# WorldGUI: Dynamic Testing for Comprehensive Desktop GUI Automation

### v2

@@ D VSCode, File Explorer Word, Adobe Acrobat
(5) Data Statistic. WorldGUI comprises 315 GUI tasks sourced from 10 widely used Windows applications, including PowerPoint, Excel, Word, Adobe Acrobat, VSCode, File Explorer, Settings, Web, Youtube, and Stable Diffusion. Here we provide the details about the task activities and project file for each application in Tab. 8. We collected 107 meta tasks from these applications, each of which was augmented 0 to 3 times based on its specific content, resulting in 208 augmented tasks. Fig. 10 presents the task distribution, provides query examples, and illustrates the distribution of meta and augmented tasks across different applications.
Applications All Task Task Activities Project File Type
PowerPoint 34 Change the content style and layout; Design new effects. project.pptx
Word 31 Formatting the content style and layout project.docx
Excel 35 Table formatting; Data management and processing project.xlsx
Adobe Acrobat 36 Automatic add electric signature; Document management project.pdf
VSCode 33 Code editing; Software configuration vscode.exe
Settings 39 Advanced personalized and safety settings; ms-Settings
File Explorer 26 File management: Add, delete, rename, and move files. explorer.exe
Web 32 Web operation web browser + URL
Youtube 28 Video and account configurations web browser + UR
Stable Diffusion 21 Media creation with complicated effects Local Run Platform
Table 8: This table shows all tasks, task activities, and project file of the desktop applications used in WorldGUI.
Figure 10: Distribution of collect tasks, selected queries, and task amount of WorldGUI. We have gathered tasks across 10 desktop applications, focusing on the use of productivity software as well as fundamental computer operations and settings.
Appendix D GUI-Thinker Reasoning Loop Algorithm
@@
Algorithm 1 GUI-Thinker Reasoning Loop Algorithm
@@
Appendix E Qualitative Results
(1) In Fig. 12, we present a successful prediction example, demonstrating that the WorldGUI can effectively plan each step for a task, accurately perceive specific elements in the GUI, and convert them into the correct action code. Additionally, we display the parsed GUI elements, which can accurately identify most content, including small icons and dense text elements.
Figure 12: We show one successful prediction of our GUI-Thinker.
@@ Appendix E Qualitative Results
(2) We provide the visualization results of using Planner-Critic, Step-Check, and Actor-Critic in Fig. 14, Fig. 15, and Fig. 16. These qualitative results demonstrate the effectiveness of these critical modules in GUI automation.
@@
(3) We also highlight some common errors encountered. 1) The model has the difficulty of obtaining the desired information when we augment the task by invoking the dropdown menu of the Settings application. As shown in the left of Fig. 17, when we click on the ’System’ button on the left, it is challenging for our model to extract the button’s position as it is hidden. Such cases require the model to have a higher level of ability to delete the content in the input box or click on the blank area. 2) As shown in the right of Fig. 17, the model has difficulty dragging a bar to achieve the desired value. 3) The model struggles with the visual choice when there is no text information in the screenshot, as shown on the left of Fig. 18. The subtask aims to select the center button, but the current model makes it hard to detect the center choice only from the screenshot. 4) The model cannot successfully locate the position of the input box, as the GUI parser will easily locate the text location ’Replace with’, it always outputs the action like clicking on the ’Replace with’, which will destroy the whole task success.

### v3

@@ D VSCode, File Explorer Word, Adobe Acrobat
Appendix C Data Statistics
Augmentation tasks type analysis. As we summarize, the real GUI scenarios include: (1) The software interface is not in its default state. (2) The human-computer interactions may start from the intermediate state of a specific task. We propose to create the augmented tasks for each meta task to stimulate authentic GUI interactions. Our augmentations lie in two main groups: (1) stimulating the intermediate states and (2) introducing diverse initial states. We divide the two groups into three exact types: Add-step, Trim-step, and Adjust-step. For Add-step, it represents various unrelated state augmentations to stimulate the scenario that we may start the agent-computer interactions in another unrelated task or interfaces, the agent should replan the task to add necessary steps. For Trim-step, it represents that we finish several steps of a long task and make the task in an intermediate state. For Adjust-step, it is usually a small modification of the existing state, such as changing the interface by clicking another Tab or clicking a button to open an unrelated dropdown menu. Most of the time, it would not require new steps to return to the target task progress. This augmentation may mislead the agent in state understanding, making them jump or miss the key steps. As shown in Figure 11, the manually created augmentations mainly belong to the add-step. Adjust-step could be the second-largest application, except for the File Explorer. Due to the low complexity of the interfaces of File Explorer, we cannot create many augmentations for adjust-step.
Task difficulty analysis. Figure 12 shows the distribution of the task difficulty across desktop applications. We annotated the task difficulty level based the subjective software usage experience. The results indicate that the tasks in Adobe Acrobat and VLC player are more challenging. The tasks in Excel, PowerPoint, and Word are more at the medium and simple levels. By considering the Success Rate and task length on these tasks, one can know that the tasks are easy for humans, but hard for current GUI agents. According to VSCode and File Explorer, and YouTube applications, the tasks are easier than in other applications. Overall, the task difficulty of created data is diverse across different applications, and there is still a need for stronger agents focusing on handling desktop-oriented GUI tasks.
The task details about the user query and the pre-actions are included in the metadata JSON file in the supplementary materials. The project file, instruction video, and augmentation files can be found in the provided data link.
Figure 10: Distribution of Software taxonomy and the distribution of task length.
Figure 11: The distribution of different augmentation types.
Figure 12: The distribution of different task difficulty.
Table 10: Task statistics of WorldGUI.
Category Software Meta Count Aug Count Meta-Task Length (Avg.)
Office Adobe Acrobat 11 55 7.1
Office Excel 12 58 4.1
Office PowerPoint 11 53 5.1
Office Word 11 52 5.1
Coding VSCode 11 45 4.4
Windows Usage File Explorer 10 34 7.1
Windows Usage Settings 14 55 5.1
Web Usage Web Browser 10 49 7.8
Web Usage YouTube 11 50 4.8
Media VLC Player 10 49 10.8
Total (Average) 111 500 6.0
Appendix D Detailed Experimental Results
Table 11 shows the detailed results of WorldGUI-Agent across individual applications in WindowsAgentArena [4] benchmark. The results of this related Windows-centric interactive GUI benchmark indicate that current the desktop GUI tasks are more challenging than web tasks. As we complete 11 out of 17 tasks in Web Browsing, a similar phenomenon is also discovered in Table 4.
Table 11: Detailed experimental results of WorldGUI-Agent across individual applications in WindowsAgentArena [4].
Domain Application #Tasks #Successes SR (%)
Web Browsing chrome 17 11 64.71
Windows Utilities clock 4 2 50.00
Windows System file_explorer 19 7 36.84
Office libreoffice_calc 24 1 4.17
Office libreoffice_writer 19 2 10.53
Windows Utilities microsoft_paint 3 1 33.33
Web Browsing msedge 13 5 38.46
Windows Utilities notepad 2 1 50.00
Windows System settings 5 4 80.00
Media & Video vlc 21 6 28.57
Coding vs_code 24 8 33.33
Windows Utilities windows_calc 3 0 0.00
Overall 154 48 31.17
Appendix E Computational Costs Discussion
The average number of execution steps and tokens consumed are shown below Table 12. The execution steps are calculated based on our experimental log files, while the token costs are sampled from representative tasks in each category by taking Actor module as an example.
Table 12: Average execution steps and token costs on different software.
Application category Average execution steps Input tokens per step (Actor) per task Output tokens per step (Actor) per task
Office $\sim$23 2350 212
Win. Usage $\sim$20 1929 108
Web $\sim$17 1637 84
Take a Windows Setting task as an example, we provide detailed time costs across different modules tested on a Laptop with AMD Ryzen 7 4800H CPU in Table 13. Task length: 6 (Generated by Planner+Planner-Critic). Total steps: 18, total time: 335.84s, total agent execution time: 412.37s. Since desktop GUI is in its early stages, computational costs are currently unavoidable. Even OpenAI’s Deep Research takes over 10 minutes in daily use. According to the post from OpenAI’s Operator, achieving 38.1% on OS-World requires over 100 steps, which is costly as well. In summary, it is still a tradeoff between performance and time costs in GUI automation.
Table 13: Time costs of each module.
Subtask Index Executed Modules Time (seconds)
0 Planner+Planner-Critic 12.53
1 Step-Check+Actor 21.23
1 Actor-Critic 13.07
2 Step-Check+Actor 17.81
2 Actor-Critic 11.86
3 Step-Check+Actor 18.43
3 Actor-Critic 11.72
4 Step-Check+Actor 22.49
4 Actor-Critic 25.31
4 Actor-Critic 10.82
5 Step-Check+Actor 17.01
5 Actor-Critic 22.36
5 Actor-Critic 26.85
5 Actor-Critic 18.99
6 Step-Check+Actor 21.79
6 Actor-Critic 23.37
6 Actor-Critic 21.12
6 Actor-Critic 19.08
Appendix F Examples of Augmentations
In this section, we present several augmentation examples in Figures 13, 14, 15, 16, 17, 18, 19. It is noted that our augmentations are not only making the first step changing but also require the agent add new step in its second step. For instance, in Figure 13, our augmentation is about click on Data tab in the ribbon, in the default software state, the Merge & Center button exhibit in the Home tab, there is no need to click on Home tab, after our augmentations, the agent should add a new task “Click on Home Tab” before it click on the Merge & Center button. Similarly, in Figure 18, the text editing buttons are under the Home Tab, if we augment the initial state with other Tab like Animation Tab, after the first step “Select the text ’US SUBMARINE DAY’ ”, the agent should add a new step like “Click on Home Tab” back to the default state for task execution. Except for adding new steps, we also present an example about adjust the step in Figure 14, as the target is about merging cells A1 to K1, we augment the initial state by selecting A2 to K2. Such a slight difference may mislead the agent to perceive such a minor difference, and the agent may jump the first step about selecting the correct cells lead to finally unsucess. In Figure 15 and Figure 19, we show two examples of introducing pop-up window in the initial state which require the agents accurately identify the pop-up windows and correctly close it by replanning the task based on the visual screenshot not only strictly planning based on inherited knowledge or the instructinal videos. In Figure 16, we show an example of changing the interface by clicking the Data tab to hide the Merge & Center button under the Home tab. In Figure 17, we complete the first step about selecting A1 to K1, which requires the agent to jump this step to reduce the time costs.
Figure 13: Augmented example of an Excel Task.
Figure 14: Augmented example of an Excel Task.
Figure 15: Augmented example of an Excel Task.
Figure 16: Augmented example of an Excel Task.
Figure 17: Augmented example of an Excel Task.
Figure 18: Augmented example of a PowerPoint Task.
Figure 19: Augmented example of a PowerPoint Task.
Appendix G WorldGUI-Agent Reasoning Loop Algorithm
@@
Algorithm 1 WorldGUI-Agent Reasoning Loop Algorithm
@@
Appendix H Qualitative Results
(1) In Figure 20, we present a successful prediction example, demonstrating that the WorldGUI can effectively plan each step for a task, accurately perceive specific elements in the GUI, and convert them into the correct action code. Additionally, we display the parsed GUI elements, which can accurately identify most content, including small icons and dense text elements.
Figure 20: We show one successful prediction of our WorldGUI-Agent.
@@ Appendix H Qualitative Results
(2) We provide the visualization results of using Planner-Critic, Step-Check, and Actor-Critic in Figure 22, Figure 23, and Figure 24. These qualitative results demonstrate the effectiveness of these critical modules in GUI automation.
@@
(3) We also highlight some common errors encountered. 1) The model has the difficulty of obtaining the desired information when we augment the task by invoking the dropdown menu of the Settings application. As shown in the left of Figure 25, when we click on the ’System’ button on the left, it is challenging for our model to extract the button’s position as it is hidden. Such cases require the model to have a higher level of ability to delete the content in the input box or click on the blank area. 2) As shown in the right of Figure 25, the model has difficulty dragging a bar to achieve the desired value. 3) The model struggles with the visual choice when there is no text information in the screenshot, as shown on the left of Figure 26. The subtask aims to select the center button, but the current model makes it hard to detect the center choice only from the screenshot. 4) The model cannot successfully locate the position of the input box, as the GUI parser will easily locate the text location ’Replace with’, it always outputs the action like clicking on the ’Replace with’, which will destroy the whole task success.